OpenStudy (hansgui):

Hello, can you help me? I have this problem: Calculate the sum of n/(2^n) from 1 to infinity, by the derivative of this serie: the sum of (x)^n from 1 to infinity. I tried, doing the derivative but i can't find the exact relation between those series, doing it in wolfram, it is 2.

OpenStudy (holsteremission):

If \(|x|<1\), then\[f(x)=\sum_{n\ge0}x^n=\frac{1}{1-x}\implies f'(x)=\sum_{n\ge1}nx^{n-1}=-\frac{1}{(1-x)^2}\]

OpenStudy (holsteremission):

You'll get the value for your series by setting \(x=\dfrac{1}{2}\).

OpenStudy (hansgui):

Thanks man (: